Subtract the following numbers in the indicated bases (do not make any conversion to a different base):

a. 46078 - 17058

b. F3216 - A6516 

c. 1100102 - 0101012

d. 12345 - 4325 


1
Expert's answer
2022-05-03T03:30:12-0400

Complement (7's) to the 1705 is 6072 (7-5=2, 7-0=7, 7-7=0,7-1=6)


Then 8' s complement to 1705 is 6072+1=6073.


Then 6073+ 4607=12702


The first bit is not used, so


Answer: 2702


b. Complement 15's to the A65 is (15-5=10=A, 15-6=9, 15-10=5) 59A.


Then 16's complement is 59A+1=59B


Then 59B+F32=14CD


The first bit is not used, so


Answer: 4CD


c.(-010101)=101010+1=101011

Then 110010+101011=1011101


The first bit is not used, so


Answer: 011101


d. Complement (4's) to the 432 is (4-2=2, 4-3=1,4-4=0) 012.

Then 5's complement is 012+1=013


Then 013+1234=1302. The first bit is not used, so


Answer: 302